A batch of n = 50 items contains m = 10 defective items. Suppose k = 10 items are selected at random and tested. How many items, k, do we need to sample, in order to get at least one defective item, with probability greater than 0.5?
Answer #1

Probability of defective = p =10/50=0.2

P( at least one )>0.5

1-P(none) >0.5

P( none defective)<0.5

0.8^x<0.5

x= ln(0.5)/ln(0.8) = 3.16

Atleast 4 items must be sampled.